# A model-agnostic gravitational-wave background characterization algorithm

### v1

@@ Appendix B Bayes factors and odds ratios
Figure 11: Results for the number of knots corresponding to the posterior recovery envelopes shown in Fig. 10. We show a heat map of the number of knots per posterior for linear (left), cubic spline (middle), and Akima spline (right) interpolation. Linear interpolation requires a greater number of knots to fit a smooth curve while the cubic and Akima splines require predominantly between 2 and 3 knots for each posterior.
@@
In this Appendix, we compare the different types of interpolation used in this work: linear, cubic spline, and Akima. We refer back to Section IV.1 for specifics on each interpolation type.
We begin by comparing the recoveries of an injected smoothed BPL signal (see Section V.3). Figure 10 presents the 95% recovery envelopes for various observation times for linear (left), cubic spline (middle), and Akima spline (right) interpolation. All three approaches can resolve the low-frequency segment of the signal well. However, of the three options, cubic spline interpolation resolves higher frequencies best where there is less detector sensitivity. Therefore, cubic splines are useful if we choose to probe all frequency regimes of a signal recovery.
The other two interpolation approaches also have their advantages. Linear interpolation will better discern whether there is strict PL behavior. Section V.2 explores an example of linear interpolation, where we consider a background subtraction signal. On the other hand, Akima spline interpolation seems to recover both PL and non-PL behavior. For small observation times (blue, orange), the 95% confidence posterior envelope generally appears to recover a PL. As the observation time increases, the Akima spline interpolation yields a smooth turnover that matches the injected signal. The limitations to using Akima spline interpolation are that there are broader posterior envelopes than for cubic spline interpolation for the BPL. However, this is likely due to cubic splines being able to describe the specific injected model with fewer parameters because a BPL is well described by cubic splines.
We now consider the number of knots required by each approach to recover the injected signal as a function of observation time, shown in Fig. 11. Linear interpolation requires the greatest number of knots for all observation times (left). This makes sense because we are trying to fit a smooth and differentiable injection with linear line segments. The cubic (middle) and Akima (right) spline interpolation methods both require mostly 2 or 3 knots in their posteriors. This means fewer knots and a more consistent distribution of the number of knots than in the linear interpolation approach.
Finally, we calculate a statistic to measure the precision with which each fitting method reconstructs the injected signal. Motivated by LVK matched-filtering searches, we utilize the mismatch [124]:
@@
[…]
@@
We calculate the mismatches between the boundaries of the 95% confidence envelopes and the average posterior, and plot the results in Fig. 12. Right away, we notice that the mismatches decrease as a function of the observation time for each of the interpolation methods. The mismatches for all three interpolation methods improve by almost three orders of magnitude between 1 week of observing time and 10 years of observing time. This trend indicates that the fits for each of the interpolation methods are improving as we increase sensitivity through longer observing times, which is what we expected.
@@
In the specific instance of the FOPT smoothed BPL, cubic splines are the optimal interpolation choice. Therefore, we employ cubic splines in Section V.3 of the main text. In contrast, linear splines are more useful for CBC-informed injection (Section V.1) and subtraction injection (Section V.2). By implementing various interpolation techniques within the main body of the paper, we highlight the respective advantages of each method, which are further elaborated upon in this appendix.

### v2

@@ Appendix C LINEAR VS SPLINE AND AKIMA INTERPOLATION
Figure 12: Results for the number of knots corresponding to the posterior recovery envelopes shown in Fig. 11. We show a heat map of the number of knots per posterior for linear (left), cubic spline (middle), and Akima spline (right) interpolation. Linear interpolation requires a greater number of knots to fit a smooth curve while the cubic and Akima splines require predominantly between two and three knots for each posterior.
@@
In this appendix, we compare the different types of interpolation used in this work: linear, cubic spline, and Akima. We refer back to Sec. IV.1 for specifics on each interpolation type.
We begin by comparing the recoveries of an injected smoothed BPL signal (see Sec. V.3). Figure 11 presents the 95% recovery envelopes for various observation times for linear (left), cubic spline (middle), and Akima spline (right) interpolation. All three approaches can resolve the low-frequency segment of the signal well. However, of the three options, cubic spline interpolation resolves higher frequencies best where there is less detector sensitivity. Therefore, cubic splines are useful if we choose to probe all frequency regimes of a signal recovery.
The other two interpolation approaches also have their advantages. Linear interpolation will better discern whether there is strict PL behavior. Section V.2 explores an example of linear interpolation, where we consider a background subtraction signal. On the other hand, Akima spline interpolation seems to recover both PL and non-PL behavior. For small observation times (blue and orange), the 95% confidence posterior envelope generally appears to recover a PL. As the observation time increases, the Akima spline interpolation yields a smooth turnover that matches the injected signal. The limitations to using Akima spline interpolation are that there are broader posterior envelopes than for cubic spline interpolation for the BPL. However, this is likely due to cubic splines being able to describe the specific injected model with fewer parameters because a BPL is well described by cubic splines.
We now consider the number of knots required by each approach to recover the injected signal as a function of observation time, shown in Fig. 12. Linear interpolation requires the greatest number of knots for all observation times (left). This makes sense because we are trying to fit a smooth and differentiable injection with linear line segments. The cubic (middle) and Akima (right) spline interpolation methods both require mostly two or three knots in their posteriors. This means fewer knots and a more consistent distribution of the number of knots than in the linear interpolation approach.
Finally, we calculate a statistic to measure the precision with which each fitting method reconstructs the injected signal. Motivated by LVK matched-filtering searches, we utilize the mismatch [Blackman_2017]
@@
$\langle\Omega_{\mathcal{M}},\hat{C}\rangle=4{\rm Re}\int^{\infty}_{-\infty}\frac{\Omega_{\mathcal{M}}(f)\hat{C}^{*}(f)}{\sigma^{2}(f)}df\,.$ (38)
@@
We calculate the mismatches between the boundaries of the 95% confidence envelopes and the average posterior and plot the results in Fig. 13. Right away, we notice that the mismatches decrease as a function of the observation time for each of the interpolation methods. The mismatches for all three interpolation methods improve by almost 3 orders of magnitude between 1 week of observing time and 10 years of observing time. This trend indicates that the fits for each of the interpolation methods are improving as we increase sensitivity through longer observing times, which is what we expected.
@@
In the specific instance of the FOPT smoothed BPL, cubic splines are the optimal interpolation choice. Therefore, we employ cubic splines in Sec. V.3 of the main text. In contrast, linear splines are more useful for CBC-informed injection (Sec. V.1) and subtraction injection (Sec. V.2). By implementing various interpolation techniques within the main body of the paper, we highlight the respective advantages of each method, which are further elaborated upon in this appendix.